OpenStudy (anonymous):

Help please!...... You need 770 mL of a 85% alcohol solution. On hand, you have a 45% alcohol mixture. How much of the 45% alcohol mixture and pure alcohol will you need to obtain the desired solution?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i need help with coming up with an equation :(

OpenStudy (anonymous):

o i see

OpenStudy (anonymous):

dont you minus?

OpenStudy (anonymous):

I honestly have no idea /:

OpenStudy (anonymous):

Let x and y be the volume of the 45% alcohol solution and the pure alcohol solution respectively. Solve the following equations for x and y. The first equation says that the total of the alcohol volumes on the LHS has to equal the alcohol volume on the RHS. The second equation relates the volume of the liquids on the LHS to the liquid volume on the RHS.\[\{0.45x+1.0y=0.85*770,x+y\text{ =}770\}\]\[\{x=210 \text{ml},y=560 \text{ml}\} \]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

You are welcome. As an aside, most of the "mixture" problems as I call them are solved with the same approach.

OpenStudy (anonymous):

ill keep that in mind, i have a ton to do still :(